A plane flies 500 miles with the wind and 340 miles against the wind in the same length of time. If
the speed of the wind is 20 mph, what is the speed of the plane in still air?

Answer by jorel1380(2518)

500/s+20=340/s-20
340(s+20)=500(s-20)
340s+6800=500s-10000
16800=160s
s=105
The still air speed of the plane is 105 mph..